Chapter 30, Problem 037 A long solenoid has a diameter of 11.1 cm. When a current i exists in its windings, a uniform magnetic field of magnitude B = 41.9 mT is produced in its interior. By decreasing i, the field is caused to decrease at the rate of 6.61 mT/s. Calculate the magnitude of the induced electric field (a) 1.31 cm and (b) 8.52 cm from the axis of the solenoid.
Answer #1

Solution)

Part a)

We know, induced electric field, E=0.5*(delta B/delta t)*r

E=0.5*(6.61*10^-3 T/s)*0.0131 m= 4.32*10^-5 V/m

=========

Part b)

E=0.5*(delta B/delta t)* R^2/r

E=0.5*(6.61*10^-3 T/s)*0.111^2/0.0852

E=0.000477 V/m

===============